OpenStudy (saylilbaby):

I have the answers jst need help working them out I will give a medal and fan whoever helps...

OpenStudy (saylilbaby):

(-5x^2-2)+(3x^2+7)=-2x^2+5 (y^2+4-2y^3)-(3y^2-7y+4y^3)=-2y^2-6y^3+4+7y (a+5)(a-3)=a^2+2a-15 (4p-3)^3=16p^2-24p+9 (7+x)(x^2+6x-8)=13x^2+34x-56+x^3

OpenStudy (saylilbaby):

@TheSmartOne @campbell_st @sleepyjess @texaschic101 @amistre64

OpenStudy (campbell_st):

1st one is correct 2nd one has an error in y^3

OpenStudy (campbell_st):

the 2nd one also has an error in the sign of 7y

OpenStudy (campbell_st):

3rd one is correct

OpenStudy (texaschic101):

hint : when you have a + in the middle. you do not need parenthesis -5x^2 - 2 + 3x^2 + 7 --- combine like terms -5x^2 + 3x^2 = -2x^2 -2 + 7 = 5 now put them together.. -2x^2 + 5 =========== if there is a - in between, do this.. y^2 + 4 - 2y^3 - (3y^2 - 7y + 4y^3) --- distribute the - through the parenthesis y^2 + 4 - 2y^3 - 3y^2 + 7y - 4y^3 now combine like terms -2y^3 - 4y^3 = -6y^3 y^2 - 3y^2 = -2y^2 7y -- there are no like terms, but do not forget about it 4 -- no like terms now put it all together.. -6y^3 - 2y^2 + 7y + 4 ============ its that simple...just a matter of combining like terms

OpenStudy (campbell_st):

and the last 1 looks ok... normally you should write then with the term having the highest power 1st... then in descending order of powers

OpenStudy (texaschic101):

(a + 5)(a - 3) --- solve using FOIL (first,outside,inside,last) F -- multiply first terms in each set....a * a = a^2 O -- multiply outside terms in each set...a * -3 = -3a I -- multiply inside terms on each set ... 5 * a = 5a L -- multiply last terms in each set...5 * -3 = -15 now put it all together.. a^2 - 3a + 5a - 15 -- combine like terms a^2 + 2a - 15

OpenStudy (saylilbaby):

@texaschic101 great job but can u do the (7+x)(x62+6x-8)